• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Yi Qiang

    Log entry time 16:04:00 on January25,2008

    Entry number 227102

    keyword=Day Shift Summary

    Shift Workers
    ----------------------------------------------------------
    SL: Yi Qiang
    TO: S. Tharanga Jinasundera
    
    Shift Summary
    ----------------------------------------------------------
    08:00-09:00	Production of -500 MeV Pmiss setting.
    09:00-10:50	Controlled access for spot move and lumi.
    11:00-11:40	Beam is back however the spot size is not 
    		good so we asked MCC to tune the beam
    11:40-13:52	Finished production of -500 MeV Pmiss 
    		setting.
    13:52-14:16	Moved Right HRS to 48.86 degrees.
    14:16-15:16	Finished production of -100 MeV Pmiss 
    		setting
    15:16-15:30	Moved Right HRS to 54.59 degrees.
    15:30-15:48	Reboot TS11
    15:48-16:00	Production of 0 MeV Pmiss setting
    
    
    Run Summary
    ----------------------------------------------------------
    Run#	From	to	I(uA)	Event#	Target	Comments
    
    -500 MeV production
    
    3291	07:40	08:20	60	500k	Thin Pb
    3292	08:21	08:56	60	412k	Thin Pb
    3293-3294	Cosmic
    3295-3300	Runs for beam tunning and raster check
    3301	11:40	11:48	20-60   113k    Thin Pb   
    3302	11:49	12:17	60	500k	Thin Pb
    3303	12:18	12:45	60	500k	Thin Pb
    3304	12:46	13:14	60	500k	Thin Pb
    3305	13:15	13:43	60	500k	Thin Pb
    3306	13:43	13:52	60	145k	Thin Pb
    
    -100 MeV production
    
    3307	14:16	15:16	60	439k	Thin Pb
    
    0 MeV production
    
    3308		Failed due to TS11 connection loss
    3309	15:48	-       60	-	Thin Pb